OpenStudy (anonymous):

Solve for "a" in the following: S = ut + 1/2 at^2

OpenStudy (anonymous):

Do it one step at a time. S-ut=1/2(at^2) 2(S-ut)=at^2 2(S-ut)/(t^2)=a \[a=2(s-ut)/(t ^{2})\]